# Plurals: A System for Guiding LLMs Via Simulated Social Ensembles

### v5

@@ 2.3. Multi-Agent Systems for AI Alignment
Multi-agent systems have a long history in artificial intelligence (oliveira_multi-agent_1999, ; van_der_hoek_chapter_2008, ). Now there is substantial interest in multi-agent LLM systems (pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ; hua_assistive_2024, ; tsao_multi-agent_2023, ; ni_mechagents_2024, ; khan_debating_2024, ). Our system incorporates aspects of these systems such as debate (irving_ai_2018, ) and the idea of role-based communication (pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ; zhu_role-based_2008, ).
Like our system, several multi-agent systems are explicitly designed with the goal of alignment (irving_ai_2018, ; pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ). Broadly, these systems typically center interactions between agents or agent roles. For example, several projects have explored the role of AI alignment through debate( (irving_ai_2018, ; khan_debating_2024, )). Khan et al. (khan_debating_2024, ) propose a system where a weaker model picks an answer after seeing stronger models debate, and this process of debate leads to more truthful answers for the weaker model. Other multi-agent systems center agent roles (pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ; zhu_role-based_2008, ; vahidov_pluralistic_2004, )—the idea being that agents playing distinct parts can aid human decision-makers (vahidov_pluralistic_2004, ).
@@ 3.1. Interactional Pluralism
Sorensen et al.’s typology of pluralistic AI systems is a useful starting point (sorensen_roadmap_2024, ). They distinguish between models that (1) present a spectrum of reasonable responses, (2) can be steered to reflect certain perspectives, and (3) are well-calibrated to a given population. The ability to craft custom personas aligns with the second type and our use of government datasets like ANES to generate nationally representative personas aligns with the third type.
Plurals extends this typology by allowing users to define the rules of engagement between agents: Structures shape the dynamics of information sharing and aggregation; Combination instructions provide an additional layer of control over how agents should incorporate each other’s views. This architectural pluralism is distinct from just having a plurality of agent-level views. Interactionally pluralistic AI systems enable users to control the “rules of engagement” that govern how Agents with differing profiles may deliberate. Plurals enables an architectural pluralism that is distinct from the conceptions of pluralism in Sorensen et al. (sorensen_roadmap_2024, ).
@@ 3.2. Modularity
Apart from the practical utility, this separation between Agents and Structures aligns with well-established social science frameworks. This conceptualization is most explicitly articulated in Structuration Theory by Anthony Giddens (giddens_constitution_1986, ), which explores the interplay between “agents” and larger “structures” agents exist in. Giddens aimed to transcend theories of behavior that centered exclusively on either individuals or societal structures. Similar distinctions appear across disciplines: individuals and environments in development psychology (radke-yarrow_individual_1991, ), person and situation in social psychology (furr_persons_2021, ), individual and field in sociology (grenfell_field_2012, ), and agent and environment in artificial intelligence (russell_artificial_1995, ). By using Agents and Structures as core abstractions [^2] We chose the terms “Agents” and “Structures” based on AI terminology and definitional precision. We adopt conventional AI terminology, where agents refer to autonomous entities capable of perceiving and acting within an environment (russell_artificial_1995, ). We opted for “structures” instead of “environment” to more accurately reflect our system. Structures specifically define information-sharing patterns and interaction protocols between agents, describing a more bounded space than “environment.” Unlike “environment,” which often implies indeterminacy or randomness, “structure” connotes intentional arrangement. The Cambridge Dictionary defines structure as “the way in which the parts of a system or object are arranged or organized, or a system arranged in this way” (dictionary_structure_2024, ). This definition aligns with a user-tuned abstraction., we create a modularity that resonates with different disciplines.
@@ 3.3. Grounded in Deliberation Practice
As described in subsection 2.1, our abstractions (Agents, Structures, Moderators) map to the practice of deliberation. Ryfe (ryfe_does_2005, ) breaks deliberation into (1) The organization of the encounter, (2) the deliberation within the encounter, and (3) the final product. These map onto Agent initialization (Phase 1), Structures and combination instructions (Phase 2), and moderation (Phase 3). By mirroring the components of deliberation, we ground our system in it. Of course, the utility of these abstractions in simulated agent space is less clear than with humans. However, incorporating these foundations can help build realistic simulations and test whether strategies developed in the literature can be used to improve LLM outputs.
The addition of Moderators provides practical benefits. Just as in human deliberation, it is helpful to have some summary of what transpired. In many multi-agent systems, one Agent aggregates the communications of others (hua_assistive_2024, ; tsao_multi-agent_2023, ; zhu_role-based_2008, ; chitsaz_multi-agent_2009, ). The motivation for adding auto-moderators—a feature where Moderators come up with their own moderation instructions based on the task—is based on the paradigm of “auto-prompting” in DSPy (khan_debating_2024, ).
@@ 3.4. Balancing Autonomy and Usability
Our system offers users substantial autonomy. First, we ensured that Agents can be used outside of Structures so users are not wedded to Structures. Second, both Agents and Structures are highly customizable. Agents can (as some examples) be over 100 LLMs, integrate with ANES, contain a different task than other Agents in a Structure, have custom combination instructions, different model parameters, etc. Likewise, Structures span a range of use cases and information-sharing protocols (e.g.: debates, ensembles, graphs) and have tuneable parameters. Advanced users can create their own Structures.
But we tried to balance this autonomy with usability. At a high level, we tried to build abstractions that are intuitive to use. Figure 2 shows code snippets. As discussed in subsection 3.2 and subsection 3.3, our core abstractions (Agents, Structures, Moderators) correspond to both common terminology and the basic process of deliberation. At a documentation level, our repository provides extensive examples of how to use each component. At an instantiation level, we made the decision for most of the package to be usable with very few custom arguments, leveraging defaults and templates. The drawback of defaults is that “artifacts have politics” (winner_artifacts_1980, ), and so this imposes certain principles on users. For example, many of the templates (apart from debate) are deliberative rather than agonistic—emphasizing building on outputs rather than arguing. By extracting our default templates to a single human-readable file on GitHub, we make these defaults more legible to users—providing some informational autonomy.
@@ 4.1.1. Component Description
  • Combination Instructions: Combination instructions define how Agents combine information from other Agents to complete the task. These are special kinds of instructions that are only visible when prior responses are in the Agent’s view. Users can rely on templates or create their own. We provide, and empirically test, templates inspired by deliberative democracy—spanning first-wave (reason-giving) and second-wave (perspective-valuing) deliberation ideals (bachtiger_deliberative_2018, ). Other templates include (e.g.) a “critique and revise” template based on Constitutional AI (bai_constitutional_2022, ) and a template inspired by New York state’s juror deliberation instructions (new_york_state_criminal_2024, ).
  • Knowledge: Conceptually, Agents differ in the knowledge that they have. Currently, we rely on the ability to use different models as a way to leverage distinct knowledge. Different models likely differ in training data and human refinement, leading to divergent priors (ashkinaze_seeing_2024, ). Users can also use retrieval-augmented generation (RAG) libraries with our system. For example, users can retrieve relevant documents for a task and add these to an Agent’s system instructions. We plan on adding more native support for RAG in future iterations of the system.
@@
(b) In a moderated ensemble, nationally representative Agents brainstorm ways to encourage recycling. Then a moderator with a persona inspired by divergent creativity literature (ashkinaze_how_2024, ) summarizes responses with custom combination instructions.
@@
Figure 2. Plurals allows users to create complex and customizable deliberations with a few lines of intuitive code. These code snippets are annotated with the features they display. For up-to-date usage, see the GitHub repository and associated documentation.
@@ 4.1.2. Implementation
System instructions can be instantiated directly by the user or by using our persona-based methods. When using persona-based methods, the full system instructions are a combination of a specific persona and a persona template which gives more instructions on how to enact that persona. See 2(a) for an example. In that example, there is a specific persona from ANES “You are a…” and then a template from second-wave deliberation that formats the persona. (Users can make their own persona templates, too—it is a string with a \${persona} placeholder.) The logic for bracketing out a specific persona from a persona template is to facilitate the ablation of an Agent’s identity versus additional instructions for how to apply that identity.
Specific personas can be inputted by the user (e.g: “A graphic designer”) or drawn from American National Election Studies (ANES) [^4] Specifically, we are using the ANES pilot dataset from February 2024., as in Argyle et al. (argyle_out_2023, ). When using ANES, our system finds a real individual satisfying some criteria and then creates a persona based on the totality of this individual’s attributes. Sampling is always probability-weighted, so the probability of a citizen being simulated matches their national sample probability weight. Because ANES is nationally representative, the marginal distribution of Plurals-generated personas matches that of the general population. Code snippet 2(d) (top panel), shows initializing Agents based on specific criteria (e.g: California resident below the age of 40) using the query_str method, which searches ANES through a Pandas string [^5] For accessibility we have a helper function which prints a human-readable mapping of ANES variables.. For convenience, we also support an ideology method (ideology=’liberal’) and initializing randomly selected ANES citizens (persona=’random’, 2(a)). The latter can be used to quickly draw up nationally representative “citizen assemblies” (2(b)).
ANES is just one possible generator of data-driven personas, and in future iterations, we aim to provide additional persona-generation methods. We chose ANES as our initial dataset for the following reasons. First, it has been used in prior work—most notably, Argyle et al. (argyle_out_2023, ). Second, ANES has data on political ideologies, supporting the core motivation of this system—testing whether LLM outputs can be improved through pluralism. Third, ANES is updated more frequently than other nationally representative datasets like the U.S. census.
@@ 4.2. Structures
Figure 3. Current Structures that Plurals supports: Chain, Graph, Debate, and Ensemble. A Chain is a sequence of agents arranged in a customizable order. It takes a list of agents with arguments: last_n (defines the number of previous responses each agent should see), cycles (determines whether to repeat deliberation), and shuffle (decides whether to reorder the agents in each cycle). A Graph (accepting a last_n argument) is a directed acyclic graph of agents where users provide agents and edges, enabling deliberation to proceed through the graph where ($A\rightarrow B$) implies B will see A’s responses. Debate involves agents engaging in back-and-forth discussions, also incorporating the cycles and last_n parameters. An Ensemble is a list of agents processing tasks in parallel, where users provide an agent and a task; this structure utilizes the cycles and last_n parameter. Plurals also supports the creation of custom structures (Appendix B).
@@ 4.2.1. Component Description
  • Amount of information shared: Chains, Debates, and DAGs have a parameter called last_n that controls how many prior responses each Agent can see. For DAGs, the density of the network can be thought of as an amount of information sharing as well. Ensembles are a basic structure where no information is shared; Agents process tasks in isolation.
@@ 4.2.2. Implementation
Existing structures we have include Chains, Graphs, Debates, and Ensembles. In an “Ensemble” no information is shared and Agents process requests in parallel. A “Chain” is a highly flexible Structure where agents build upon each other’s answers with deliberation optionally rewired on each cycle (2(d), bottom panel). There, three Agents will build on each other’s output for three cycles. The initial order is agent1->agent2->-agent3 but because shuffle=True, the order will change each cycle. Debates involve a back-and-forth between two agents (2(d), top panel).
The Graph structure enables users to create directed acyclic graphs (DAGs) of Agents, processing tasks via Kahn’s algorithm for topological ordering. DAGs allow “upweighting” certain voices by increasing their connectedness. In 2(c), Agents critique and revise a company memo using the combination_instructions = ‘critique_revise’ template. A woman ANES Agent’s output is fed forward to all of the other Agents (so they see that Agent’s responses when answering). Then a Moderator summarizes all responses.
The possibility space of potential structures is vast. Our existing structures provide a lot of customizability. But some users will want a structure that has a different behavior than what can be accomplished via existing structures. Consequently, we built the package so that advanced users can easily create their own custom structures, leveraging the polymorphic design of the structure classes (more details in Appendix B).
@@ 4.3.2. Implementation
Moderators can be useful when users want an Agent who will not participate in deliberation but merely summarize it. For example, users may want to have a chain or ensemble of liberals with an independent Moderator summarizing responses at the end. As with other components, we offer pre-defined templates for Moderators. We support various pre-defined moderator instructions such as “information aggregators” or “synthesizers”. Inspired by auto-prompting libraries such as DSPy (khattab_dspy_2023, ), we also support Auto-Moderators. Given a task, an Auto-Moderator will ask itself what the system instructions of a Moderator should be for the task it was assigned. Auto-Moderators are initialized through system_instructions=’auto’ (bottom panel of 2(d)).
Table 1. A summary of empirical case studies. Mechanistic fidelity studies support claims we make about how the system is operating. Efficacy checks compare the output of the system against zero-shot. One case study explores Plurals as a system for managing LLM abstentions. See Appendix Table 3 for multilevel logistic regressions of efficacy experiments.
Study No. Type System Component(s) Result
1 (Appendix C) Mechanistic Fidelity Personas ANES personas yield more diverse responses than single-attribute personas (100% of comparisons for Claude Sonnet, 95% of comparisons for GPT-4o).
2 (subsection 5.2) Mechanistic Fidelity Combination Instructions We developed instructions based on democratic deliberation literature. The fidelity of (a subset of) these instructions was validated by crowdworkers (89% accuracy when comparing the model’s output to the given instructions).
3 (subsection 5.3) Efficacy Personas, Ensembles, Moderators Conservatives preferred solar panel company ideas from a simulated focus group of conservatives over zero-shot generation in 88% of trials.
4 (subsection 5.4) Efficacy Personas, DAGs Liberals preferred charter school ideas from a simulated focus group of liberals over chain-of-thought zero-shot generation in 69% of trials.
5 (subsection 5.5) Efficacy Personas, DAGs Liberals preferred homeless shelter proposals from a simulated focus group of liberals over chain-of-thought zero-shot generation in 66% of trials.
6 (subsection 5.6) Moderation Moderators Using Plurals, end-users can create steerable LLM guardrails (91% accuracy in a value-based abstention experiment).

### v6

@@ 2.3. Multi-Agent Systems for AI Alignment
Multi-agent systems have a long history in artificial intelligence (Oliveira et al., 1999; van der Hoek and Wooldridge, 2008). Now there is substantial interest in multi-agent LLM systems (Pang et al., 2024; Mangal et al., 2024; Hu et al., 2024; Hua et al., 2024; Tsao, 2023; Ni and Buehler, 2024; Khan et al., 2024). Our system incorporates aspects of these systems such as debate (Irving et al., 2018) and the idea of role-based communication (Pang et al., 2024; Zhu and Zhou, 2008).
Like our system, several multi-agent systems are explicitly designed with the goal of alignment (Irving et al., 2018; Pang et al., 2024; Mangal et al., 2024; Hu et al., 2024). Broadly, these systems typically center interactions between agents or agent roles. For example, several projects have explored the role of AI alignment through debate  (Irving et al., 2018; Khan et al., 2024). Other multi-agent systems center agent roles (Pang et al., 2024; Mangal et al., 2024; Zhu and Zhou, 2008; Vahidov and Fazlollahi, 2004)—the idea being that agents playing distinct parts can aid human decision-makers (Vahidov and Fazlollahi, 2004).
@@ 3.1. Interactional Pluralism
Sorensen et al.’s typology of pluralistic AI systems is a useful starting point (Sorensen et al., 2024). They distinguish between models that (1) present a spectrum of reasonable responses, (2) can be steered to reflect certain perspectives, and (3) are well-calibrated to a given population. The ability to craft custom personas aligns with the second type and our use of government datasets like ANES to generate nationally representative personas aligns with the third type.
Plurals extends this typology by allowing users to define the rules of engagement between agents: Structures shape the dynamics of information sharing and aggregation; Combination instructions provide an additional layer of control over how agents should incorporate each other’s views. This architectural pluralism is distinct from just having a plurality of agent-level views. Interactionally pluralistic AI systems enable users to control the “rules of engagement” that govern how Agents with differing profiles may deliberate. Plurals enables an architectural pluralism that is distinct from the conceptions of pluralism in Sorensen et al. (Sorensen et al., 2024).
@@ 3.2. Modularity
Apart from the practical utility, this separation between Agents and Structures aligns with well-established social science frameworks. This conceptualization is most explicitly articulated in Structuration Theory by Anthony Giddens (Giddens, 1986), which explores the interplay between “agents” and the “structures” they exist in. Giddens aimed to transcend theories of behavior that centered exclusively on either one. Similar distinctions appear across disciplines: individuals and environments in development psychology (Radke-Yarrow, 1991), person and situation in social psychology (Furr and Funder, 2021), individual and field in sociology (Thomson, 2012), and agent and environment in artificial intelligence (Russell and Norvig, 1995). By using Agents and Structures as core abstractions [^2] We chose the terms “Agents” and “Structures” based on AI terminology and definitional precision. We adopt conventional AI terminology, where agents refer to autonomous entities capable of perceiving and acting within an environment (Russell and Norvig, 1995). We opted for “structures” instead of “environment” to more accurately reflect our system. Structures specifically define information-sharing patterns and interaction protocols between agents, describing a more bounded space than “environment.” Unlike “environment,” which often implies indeterminacy or randomness, “structure” connotes intentional arrangement. The Cambridge Dictionary defines structure as “the way in which the parts of a system or object are arranged or organized, or a system arranged in this way” (Dictionary, 2024). This definition aligns with a user-tuned abstraction., we create a modularity that resonates with different disciplines.
@@ 3.3. Grounded in Deliberation Practice
As described in subsection 2.1, our abstractions (Agents, Structures, Moderators) map to the practice of deliberation. By mirroring the components of deliberation, we ground our system in it. Of course, the utility of these abstractions in simulated agent space is less clear than with humans. However, incorporating these foundations can help build realistic simulations and test whether strategies developed in the literature can be used to improve LLM outputs.
The addition of Moderators provides practical benefits. Just as in human deliberation, it is helpful to have some summary of what transpired. In many multi-agent systems, one Agent aggregates the communications of others (Hua et al., 2024; Chitsaz and Seng, 2009). The motivation for adding auto-moderators—a feature where Moderators come up with their own moderation instructions based on the task—is based on the paradigm of “auto-prompting” in DSPy (Khattab et al., 2023).
@@ 3.4. Balancing Autonomy and Usability
Our system offers users autonomy. First, we ensured that Agents can be used outside of Structures so users are not wedded to Structures. Second, both Agents and Structures are highly customizable. Agents can (as some examples) be over 100 LLMs, integrate with ANES, contain a different task than other Agents in a Structure, have custom combination instructions, different model parameters, etc. Likewise, Structures span a range of information-sharing protocols (e.g.: debates, ensembles, graphs) and have tuneable parameters. Advanced users can create their own Structures.
But we tried to balance this autonomy with usability. First, we aimed for intuitive abstractions. Figure 2 shows code snippets of Agents, Structures, and Moderators working together. Second, we provide extensive documentation on how to use each component. Third, most of the package is usable with very few custom arguments, leveraging defaults and templates. The drawback of defaults is that “artifacts have politics” (Winner, 1980), and so this imposes certain principles on users. For example, many of the templates (apart from debate) are deliberative rather than agonistic—emphasizing building on outputs rather than arguing. By extracting our default templates to a single human-readable file on GitHub, we make these defaults more legible to users—balancing usability with informational autonomy.
@@ 4.1.1. Component Description
  • Combination Instructions: Combination instructions define how Agents combine information from other Agents to complete the task. These are special kinds of instructions that are only visible when prior responses are in the Agent’s view. Users can rely on templates or create their own. We provide, and empirically test, templates inspired by deliberative democracy—spanning first-wave (reason-giving) and second-wave (perspective-valuing) deliberation ideals (Bächtiger et al., 2018). Other templates include (e.g.) a “critique and revise” template based on Constitutional AI (Bai et al., 2022) and a template inspired by New York state’s juror deliberation instructions (State, 2024).
  • Knowledge: Conceptually, Agents differ in the knowledge that they have. Currently, we rely on the ability to use different models as a way to leverage distinct knowledge. Different models likely differ in training data and human refinement, leading to divergent priors (Ashkinaze et al., 2024a). Users can also use retrieval-augmented generation (RAG) libraries with our system. For example, users can retrieve relevant documents for a task and add these to an Agent’s system instructions. We plan on adding more support for RAG in future iterations.
@@
(b) In a moderated ensemble, nationally representative Agents brainstorm ways to encourage recycling. Then a moderator with a persona inspired by divergent creativity literature (Ashkinaze et al., 2024b) summarizes responses with custom combination instructions.
@@
Figure 2. Plurals allows users to create complex and customizable deliberations with a few lines of intuitive code. These code snippets are annotated with the features they display. For up-to-date syntax and snippets, see the GitHub repository and associated documentation.
[…]
Plurals enables different kinds of Agent initialization strategies, information-sharing protocols (Structures), and Moderator types. These code snippets show how users can: initialize Agents from nationally representative datasets, create moderated nationally representative ensembles, create critique-and-revise chains, run AI debates, and use auto-moderators.
@@ 4.1.2. Implementation
System instructions can be instantiated directly by the user or by using our persona-based methods. When using persona-based methods, the full system instructions are a combination of a specific persona and a persona template which gives more instructions on how to enact that persona. See 2(a) for an example. In that example, there is a specific persona from ANES (“You are a…”) and then a template from second-wave deliberation that formats the persona. (Users can make their own persona templates, too—it is a string with a \${persona} placeholder.) The logic for bracketing out a specific persona from a persona template is to facilitate the ablation of an Agent’s identity versus additional instructions for how to apply that identity.
Specific personas can be inputted by the user (e.g.: “A graphic designer”) or drawn from American National Election Studies (ANES) [^4] Specifically, we are using the ANES pilot dataset from February 2024., as in Argyle et al. (Argyle et al., 2023b). When using ANES, our system finds a real individual satisfying some criteria and then creates a persona based on the totality of this individual’s attributes. Sampling is always probability-weighted, so the probability of a citizen being simulated matches their national sample probability weight. Because ANES is nationally representative, the marginal distribution of Plurals-generated personas matches that of the general population. Code snippet 2(d) (top panel), shows initializing Agents based on specific criteria (e.g.: California resident below the age of 40) using the query_str method, which searches ANES through a Pandas string [^5] For accessibility we have a helper function which prints a human-readable mapping of ANES variables.. For convenience, we also support an ideology method (ideology=’liberal’) and initializing randomly selected ANES citizens (persona=’random’, 2(a)). The latter can be used to quickly draw up nationally representative “citizen assemblies” (2(b)).
ANES is just one possible generator of data-driven personas, and in future iterations, we aim to provide additional persona-generation methods. We chose ANES as our initial dataset for the following reasons. First, it has been used in prior work—most notably, Argyle et al. (Argyle et al., 2023b). Second, ANES has data on political ideologies, supporting the core motivation of this system—testing whether LLM outputs can be improved through pluralism. Third, ANES is updated more frequently than other nationally representative datasets like the U.S. census.
@@ 4.2. Structures
Figure 3. Current Structures that Plurals supports: Chain, Graph, Debate, and Ensemble. A Chain is a sequence of agents arranged in a customizable order, with the option to shuffle the order on each cycle. A Graph is a directed acyclic graph of agents where users provide agents and edges, enabling deliberation to proceed through the graph where ($A\rightarrow B$) implies B will see A’s responses. Debate involves exactly two agents engaging in back-and-forth discussions. An Ensemble is a list of agents processing tasks in parallel. Plurals also supports the creation of custom structures (Appendix A).
[…]
This conceptual diagram shows the different kind of Structures that Plurals supports. A Chain is a sequence of agents arranged in a customizable order, with the option to shuffle the order on each cycle. A Graph is a directed acyclic graph of agents where users provide agents and edges, enabling deliberation to proceed through the graph where ($A\rightarrow B$) implies B will see A’s responses. Debate involves exactly two agents engaging in back-and-forth discussions. An Ensemble is a list of agents processing tasks in parallel.
@@ 4.2.1. Component Description
  • Amount of information shared: Chains, Debates, and DAGs have a parameter called last_n that controls how many prior responses each Agent can see. For DAGs, the density of the network can be thought of as the amount of information shared. Ensembles are a basic structure where no information is shared; Agents process tasks in isolation.
@@ 4.2.2. Implementation
Existing structures we have include Chains, Graphs, Debates, and Ensembles. In an “Ensemble” no information is shared and Agents process requests in parallel. A “Chain” is a highly flexible Structure where agents build upon each other’s answers with deliberation optionally rewired on each cycle (2(d), bottom panel). There, three Agents will build on each other’s output for three cycles. The initial order is agent1->agent2->agent3 but because shuffle=True, the order will change each cycle. Debates involve a back-and-forth between two agents (2(d), top panel).
The Graph structure enables users to create directed acyclic graphs (DAGs) of Agents, processing tasks via Kahn’s algorithm for topological ordering. DAGs allow “upweighting” certain voices by increasing their connectedness. In 2(c), Agents critique and revise a company memo using the combination_instructions = ‘critique_revise’ template. A woman ANES Agent’s output is fed forward to other Agents (so they see that Agent’s responses when answering). Then a Moderator summarizes all responses.
The possibility space of potential structures is vast. Our existing structures provide a lot of customizability. But some users will want a structure that has a different behavior than what can be accomplished via existing structures. Consequently, we built the package so that advanced users can easily create their own custom structures, leveraging the polymorphic design of the structure classes (more details in Appendix A).
@@ 4.3.2. Implementation
Moderators can be useful when users want an Agent who will not participate in deliberation but merely summarize it. For example, users may want to have a chain or ensemble of liberals with an independent Moderator summarizing responses at the end. As with other components, we offer pre-defined templates for Moderators. We support various pre-defined moderator instructions such as “information aggregators” or “synthesizers”. Inspired by auto-prompting libraries such as DSPy (Khattab et al., 2023), we also support Auto-Moderators. Given a task, an Auto-Moderator will ask itself what the system instructions of a Moderator should be for the task it was assigned. Auto-Moderators are initialized through system_instructions=’auto’ (bottom panel of 2(d)).
Table 1. A summary of empirical case studies. Mechanistic fidelity studies support claims we make about how the system is operating. Efficacy checks compare the output of the system against zero-shot. One case study explores Plurals as a system for managing LLM abstentions. The leftmost column lists the study number and where to find more details. See Supplemental Materials (SM) 2 for multilevel logistic regressions of efficacy experiments.
Study No. Type System
Component(s)
Result
1 (Appendix B) Mechanistic
fidelity
Personas Using ANES personas yields more diverse responses over single-attribute personas (100% of comparisons for Claude Sonnet, 95% of comparisons for GPT-4o).
2 (Appendix C) Mechanistic
fidelity
Combination
instructions
We developed instructions based on democratic deliberation literature. The fidelity of (a subset of) these instructions was validated by crowdworkers (89% accuracy when comparing the model’s output to the given instructions).
3 (SM 3) Efficacy Personas, Ensembles, Moderators Conservatives preferred solar panel company ideas from a simulated focus group of conservatives over zero-shot generation in 88% of trials.
4 (SM 4) Efficacy Personas, DAGs Liberals preferred charter school ideas from a simulated focus group of liberals over chain-of-thought zero-shot generation in 69% of trials.
5 (SM 5) Efficacy Personas, DAGs Liberals preferred homeless shelter proposals from a simulated focus group of liberals over chain-of-thought zero-shot generation in 66% of trials.
6 (Appendix D) Moderation Moderators Using Plurals, end-users can create steerable LLM guardrails (91% accuracy in a value-based abstention experiment).